OpenStudy (loser66):

Let A be a nonzero additive abelian group. We say A is decomposable provided there exists groups B and C, both nonzero, such that \(A\simeq B\times C\). A nonzero additive abelian group that is not decomposable is said to be indecomposable. We say that A is uniform provided every nonzero subgroup of A is indecomposable. Prove that a nonzero additive abelian group is uniform if and only if the intersection of any two nonzero subgroups must also be nonzero. Please, help.

OpenStudy (loser66):

@myininaya

OpenStudy (anonymous):

why not @loser69

OpenStudy (fibonaccichick666):

ask @zepdrix ,because I do not know

myininaya (myininaya):

I have a lot of terms to look up.

myininaya (myininaya):

A is isomorphic to Cartesian product of B and C?

OpenStudy (loser66):

yup

OpenStudy (loser66):

Question: in additive abelian group, is it true that : if H1, H2, .... Hn are subgroup of G, then G = H1+H2 +...+Hn ?

OpenStudy (loser66):

when Hi is indecomposable. For example : subgroups of Z , + are A = 2Z, B = 3Z, then Z = A+B

myininaya (myininaya):

I don't know...I can't recall that theorem or even know if I know how to prove it... so we want to suppose A is an abelian nonzero group and that it is uniform...And we want to show B intersect C is nonzero given B and C are nonzero subgroups of A. So A is a uniform abelian nonzero group... This means what exactly... Well are given that B and C are indecomposable.

OpenStudy (loser66):

I got stuck at this problem because I see the conflict between the problems. I have another question like: Prove that the additive group Z is uniform. To me, that means (Z,+) is uniform, then, the intersection of any of 2 subgroups must be non zero, but obviously \(2Z\cap3Z=\{0\}\)

OpenStudy (loser66):

So, how (Z,+) is uniform?

myininaya (myininaya):

This problem makes me cry. I can't look at it anymore. I'm sorry @loser66 .

OpenStudy (loser66):

It's ok, friend. I do appreciate what you do for me.

myininaya (myininaya):

I think Zarkon could be more help on this.

myininaya (myininaya):

I honestly try to do a little research before I gave up...

myininaya (myininaya):

Just researching some of the definitions...And when I looked at all of it...It was like written in a different language.

OpenStudy (loser66):

:) Ok, let me ask others for help. Thanks a lot.

OpenStudy (anonymous):

hmm i have no idea, but i am sure that \[3\mathbb{Z}\cap 2\mathbb{Z}\neq \{0\}\]

OpenStudy (loser66):

how?

OpenStudy (loser66):

\[3\mathbb Z=\{multiple~of~3\}\] while \[2\mathbb Z=\{multiple~of~2\}\]

OpenStudy (anonymous):

6

OpenStudy (loser66):

yes! :)

OpenStudy (anonymous):

So, the intersection isn't trivial.

OpenStudy (loser66):

but we have a bunch of subgroups of Z, check each pair?

OpenStudy (anonymous):

no check \[n\mathbb{Z}\cap m\mathbb{Z}\]

OpenStudy (loser66):

where n is odd and m is even, right? general case only, right?

OpenStudy (anonymous):

still have no idea how to answer your original question

OpenStudy (anonymous):

you are thinking too hard doesn't matter what m and n are maybe you are getting them confused with \[\mathbb{Z}_n\]

OpenStudy (loser66):

nope, intersection of Z_n is always non trivial

OpenStudy (anonymous):

the answer lies in god

OpenStudy (loser66):

I got it. Thanks all